OpenStudy (anonymous):

The lengths of two sides of a triangle are shown below: Side 1: 8x2 - 5x - 2 Side 2: 7x - x2 + 3 The perimeter of the triangle is 4x3 - 3x2 + 2x - 6. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points) Can someone explain to me how to do this?

jimthompson5910 (jim_thompson5910):

What do you get when you add side1 to side2?

OpenStudy (anonymous):

7x^2 + 2x +1 ?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

that's part A

jimthompson5910 (jim_thompson5910):

since that says "Part A: What is the total length of the two sides, 1 and 2, of the triangle?"

jimthompson5910 (jim_thompson5910):

Part B: What is the length of the third side of the triangle?

OpenStudy (anonymous):

How do I find out the third side? does the perimeter do something? I think that the perimeter is all sides added, so does that mean I have to subtract the perimeter from part a?

jimthompson5910 (jim_thompson5910):

Well to get the perimeter of the triangle, you add the 3 sides

jimthompson5910 (jim_thompson5910):

Perimeter = side1 + side2 + side3

jimthompson5910 (jim_thompson5910):

So yes, but you subtract in the other order Perimeter = side1 + side2 + side3 Perimeter - (side1 + side2) = side3

OpenStudy (anonymous):

4x^3 -10x^2 - 7 ?

jimthompson5910 (jim_thompson5910):

Perimeter - (side1 + side2) = side3 (4x^3 - 3x^2 + 2x - 6) - (7x^2 + 2x +1) = side3 4x^3 - 3x^2 + 2x - 6 - 7x^2 - 2x -1 = side3 4x^3 -10x^2 - 7 = side3 side3 = 4x^3 -10x^2 - 7 so you are correct

OpenStudy (anonymous):

Thank you, Do you know what I need to do to find out part c?

jimthompson5910 (jim_thompson5910):

Do you know what they mean when they say "closed" in part C? it's a very specific term when it comes to algebra

OpenStudy (anonymous):

no I do not, thats why I am confused.

jimthompson5910 (jim_thompson5910):

when they say "closed" they mean that adding two polynomials gives you a polynomial subtracting two polynomials gives you a polynomial In general, if an operation is closed, then you won't get some other thing that's not a polynomial

jimthompson5910 (jim_thompson5910):

example with integers adding two integers ALWAYS results in another integer 2+5 = 7 8+3 = 11 9+17 = 26 etc etc pick any two integers, add them, the result is an integer. So adding integers is closed (you won't add 2 integers to get a non-integer, it's just not possible)

OpenStudy (anonymous):

so mine is closed because I got a 3 terms?

jimthompson5910 (jim_thompson5910):

so it's like an exclusive club that no one else can join once you have your members set (eg: integers), you can't go outside the club to get non-integers

jimthompson5910 (jim_thompson5910):

it's closed because you got a polynomial as a result

jimthompson5910 (jim_thompson5910):

a polynomial is simply the sum of a bunch of monomials

OpenStudy (anonymous):

so it is always closed if I am adding or subtracting integers

jimthompson5910 (jim_thompson5910):

integers was an example

OpenStudy (anonymous):

I mean polynomials?

jimthompson5910 (jim_thompson5910):

yes, adding or subtracting polynomials gives you a polynomial

jimthompson5910 (jim_thompson5910):

The basic template is Polynomial1 + Polynomial2 = Polynomial3 Polynomial1 - Polynomial2 = Polynomial3

jimthompson5910 (jim_thompson5910):

It is impossible to add two polynomials and get a non-polynomial

OpenStudy (anonymous):

Thank you for your help

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

@jim_thompson5910 so for part c it is closed correct?

jimthompson5910 (jim_thompson5910):

That is correct. Adding ANY two polynomials results in some other polynomial. Therefore, polynomial addition is a closed operation. The same can be said about polynomial subtraction.